OpenStudy (anonymous):

simplify the expression: (√9x^3) (√18x)

OpenStudy (mathmate):

Recall: \(\sqrt 3 \times \sqrt 6 = \sqrt{3\times 6}\) Try to work from there.

OpenStudy (anonymous):

i need to know how to do the first part though